# Two Types of AI Existential Risk: Decisive and Accumulative

### v2

@@ 8.2 Harmonization of ethical and existential risks
Traditionally, particularly in the Anglo-American context, there has been a tendency to treat communities focusing on ethical risks and those concerned with x-risks as distinct (Section 2). However, the accumulative AI x-risk hypotheses challenge this separation, indicating that such a dichotomy is epistemically unsound in the context of AI risks.
In addition, the distinction between decisive and accumulative AI x-risks suggests a potential realignment of AI x-safety research priorities. While investigating catastrophic failure modes associated with ASI remains crucial, there is a growing need to allocate comparable attention and resources to the exploration of more subtle, long-term existential effects of AI via ethical risks. These effects, often cumulative in nature, require thorough study to understand their full impact and implications. Neglecting the accumulative aspect of AI x-risks would be a critical oversight. Failing to recognize and investigate these gradual and accumulative impacts could leave us unprepared for the nuanced existential challenges.
Moreover, the harmonization facilitates the transfer of risk mitigation strategies across different projects. A prime example of this is the evolution of interpretable models to mitigate social AI risks into mechanistic interpretability to mitigate AI x-risks. This progression demonstrates how methodologies initially developed in one domain can be adapted and enhanced for application in another, thereby ensuring a more cohesive and efficient approach to managing AI risks.
8.3 Short-long-term risk management
In the literature, there has been an implicit assumption that it is plausible to separate short-term (ethical) and long-term (existential) AI risks. As I have suggested, upon closer examination, these lines blur, revealing a complex interplay of concerns that transcend traditional time frames. The immediate risks of advanced AI systems is not solely confined to the present but also encompasses the rapidly approaching future, indicating the interconnected nature of these risks. As such, the structure of AI risk-related problems indeed represents a complex blend of immediate, intermediate, and long-term considerations, challenging our conventional understanding of risk and time.
@@ 8.3 Short-long-term risk management
Recognising the accumulative emergence of x-risks as a result of mismanagement, accumulation, or contingency path through ethical risks allows us to plan for schedules that address simultaneously both short-term and long-term risk mitigations and emphasise the importance of ongoing monitoring and evaluation. It is interesting to explore how current AI risk management frameworks (Baryannis et al., 2019; Tabassi, 2023) adapt to the accumulative and decisive hypothesis in this paper.
8.4 Enhancing Understanding and Awareness of AI Risks
The critical link between non-existential social risks and x-risks shows the importance of elevating public awareness and education regarding the radical impacts of AI. This effort is geared towards empowering individuals and communities to participate meaningfully in conversations surrounding responsible AI development.
Decisive x-risks, characterized by their immediate and striking effects, has tended to capture public attention. They resonate with cognitive biases, tap into inherent fears, or are amplified by sensationalist media portrayals. This dynamic may result in a distorted perception of AI x-risks, where the immediacy and drama of these risks eclipse the more subtle and gradually developing accumulative x-risks. In contrast, the accumulative risks, owing to their incremental nature and less obvious long-term impacts, often do not generate the same level of public interest or urgency. They typically do not provoke the instant emotional reaction seen with decisive x-risks, potentially leading to delayed public recognition and response until these risks escalate to critical levels.
Therefore, it is important to steer the conversation on AI risks away from purely dystopian or utopian narratives. While such narratives may be engaging, they frequently lack the necessary depth and complexity for constructive discussion and effective risk management. A more nuanced and informed approach to discussing AI risks is crucial for fostering a balanced understanding and preparing for the future with AI.
8.5 Mitigating x-catastrophes by limited connectivity
Finally, the distinction between two types of AI x-risks explores the strategy of mitigating decisive AI-related x-catastrophes by deliberately avoiding universal connectivity to ASI. Understanding the risk landscapes through systems analysis reveals the potential benefits of establishing multiple, critical nodes. This strategic decentralization serves as a protective measure, enhancing global resilience against the spread of catastrophic events, particularly those instigated by ASI.
If the paper clip maximizer is not expansively connected to humanity, the potential for decisive AI x-risk substantially diminishes. Consequently, the discussion so far advocates for a cautious stance against the globalization of ASI.
@@ 9 Concluding remarks
I began by asking whether the conception of x-risk from AI systems relate to the conception of non-existential social risks. Conventional discussions of AI x-risks often respond negatively to this question, and overlook an alternative way by which AI could bring about x-catastrophes. Against this background, I formulated an alternative view of AI x-risk. I used systems analysis perspective to differentiate between the two conceptions of AI x-risks in relation to two distinct causal paths that could have AI x-catastrophes as their endpoint.
The two type of causal paths differ based on the way they instantiate three path features: non-linearity, connectedness, and dynamic system adaptation. Causal pathways that feature humanity cascade effect of ASI, super-connectedness of infrastructures to ASI, and system adaptation to unidirectional feedback loops, can end in decisive ASI x-catastrophes. Causal pathways that feature local causality with compounded systemic impact, selective-connectedness of infrastructures, and dynamic system adaptation to multidirectional feedback loops, can end in accumulative AI x-catastrophes. So x-risks should be about those events that occur in pathways and mitigation strategies targetted at them. I used a scenario termed perfect storm MISTER to propose how the accumulative model can give rise to AI x-catastrophes. I also discussed five governance implications of my proposed expanded conception of AI x-risks.
Going forward, I think there is no inherent reason to consider that the accumulative hypothesis is any less likely than the decisive view. The need to further substantiate the accumulative hypothesis is apparent. This could be effectively addressed through the development of computational simulations using a system dynamics approach, aligning with the perspective I have established in analyzing the two types of AI existential risks. Finally, my discussion has so far only differentiated between two types of AI x-risks. The possibility of additional categories remains an open question. The implications of my arguments for assessing the probabilities of catastrophic outcomes and existential risks are significant and warrant further exploration in future research.

### v3

@@ 8.2 Unifying social and existential risks
Traditionally, there has been a tendency to treat concerns focused on ethical risks and those concerned with x-risks as distinct (see Section 2). However, the accumulative AI x-risk hypotheses challenge this assumed separation, indicating that such a dichotomy is epistemically unsound in the context of AI x-risks.
The distinction between decisive and accumulative AI x-risks suggests a needed rebalancing of safety research priorities. While investigating ASI failure modes remains important, equal attention should be given to understanding how social risks compound into existential threats. These compounding effects demand systematic study to understand how they might escalate into existential threats. Focusing solely on decisive scenarios while neglecting accumulative pathways would leave us blind to critical risks that build gradually through system interactions.
Risk framework unification enables critical risk mitigation methodologies to transfer between domains. The evolution from interpretable models for social risks to mechanistic interpretability for existential risks demonstrates this potential. However, we must systematically identify what additional components are needed to scale current ethical frameworks to address catastrophic risks.
@@
Recognizing the accumulative emergence of x-risks as a result of mismanagement, accumulation, or contingency path through ethical risks allows us to plan for schedules that address simultaneously both short-term and long-term risk mitigations and emphasise the importance of ongoing monitoring and evaluation. It is interesting to explore how current AI risk management frameworks (Baryannis et al., 2019; Tabassi, 2023) adapt to the accumulative and decisive hypothesis in this paper.
@@ 9 Concluding remarks
This paper critically examined how AI existential risks are conceptualized in the current literature. The dominant framing, which I term the decisive AI x-risk hypothesis, focuses on scenarios where AGI or ASI directly causes catastrophic outcomes at a decisive moment, exemplified by thought experiments like the paperclip maximizer. Through systems analysis, I proposed and defended an alternative causal pathway to AI existential catastrophes: the accumulative AI x-risk hypothesis. This hypothesis suggests that existential risks can emerge from the compounding of multiple lower-severity disruptions that gradually weaken systemic resilience until a triggering event causes unrecoverable collapse. The perfect storm MISTER scenario illustrates this pathway, demonstrating how interactions between different types of AI risks could lead to existential catastrophe through progressive system degradation rather than just a single decisive event.
This accumulative perspective has important implications for AI risk governance. While the decisive view tends to treat x-risks as categorically distinct from other AI risks, the accumulative hypothesis reveals critical relationships between different risk types and their potential to compound. This suggests the need for integrated governance approaches that address both immediate AI risks and their potential accumulative effects. The systems analysis approach shows how seemingly manageable risks can interact and amplify through feedback loops and network effects, creating emergent threats to systemic stability that may be overlooked when risks are analyzed in isolation.
Several key questions remain from this analysis that warrant further investigation.
First, we need better methods for identifying when disruptions become critically significant. This requires developing sophisticated monitoring systems that can track not just individual AI incidents, but also their cumulative effects and interactions across different subsystems. Such monitoring systems would need to establish clear thresholds that signal when accumulating disruptions are approaching dangerous levels.
Second, while this paper introduces systems analysis to AI x-risk conceptualization and evaluation, we need more structured approaches for analyzing how risks accumulate. This includes developing formal frameworks for mapping causal chains, identifying key feedback loops, and understanding how different types of AI risks interact within complex global systems. Such frameworks would help systematically identify potential accumulative pathways to existential risk that might be overlooked by current approaches.
Third, the quantification of accumulative AI x-risks presents unique methodological challenges. Unlike decisive scenarios where single events trigger catastrophic outcomes, accumulative risks involve complex interactions over time that are harder to model mathematically. We need new methods for calculating how multiple smaller risks combine and amplify, and how to assess the probability of system-wide failures emerging from these interactions. These challenges require innovations in risk modeling that can capture the dynamic, non-linear nature of risk accumulation.
Looking ahead, there is no inherent reason to consider the accumulative hypothesis less plausible than the decisive view. Future work should focus on developing computational simulations using system dynamics to further substantiate the accumulative hypothesis and explore its implications. While this paper has contrasted decisive and accumulative pathways, other potential causal pathways to AI x-risk may exist. The framework developed here provides a foundation for investigating these possibilities through formal modeling of how different AI risks might interact and compound over time.
Acknowledgments. I thank Mazviita Chirimuuta, Iason Gabriel, Jan Matusiewicz, Mario Guenther, Matthijs Maas, and John Zerilli for their valuable feedback. I am also grateful to participants in the AI, Data & Society group at the University of Edinburgh, the Ethics Reading Group at the University of Edinburgh, the Workshop on Digital Democracy at the University of Zurich, the Sociotechnical AI Safety Workshop in Rio de Janeiro, and members of the Collective Intelligence Institute. I thank Bálint Gyevnár for helping to create the figures.